OpenStudy (anonymous):

Cheyanne is saving for a car. The amount she has saved for the car, f(x), can be represented by the function f(x) = 65x + 750, where x is the number of weeks she has saved. Suppose 750 changed to 700. What does this mean in the context of the problem? Changing 750 to 700 means the initial amount she saved decreased. Changing 750 to 700 means the initial amount she saved increased. Changing 750 to 700 means the amount she saves each week decreased. Changing 750 to 700 means the amount she saves each week increased.

OpenStudy (anonymous):

pls help 4 yung female

OpenStudy (anonymous):

nvm i think its A

OpenStudy (anonymous):

what about this one Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form. x – 4 = –1(y – 2) x – 2 = –1(y – 4) y – 2 = –1(x – 4) y – 4 = –1(x – 2)

OpenStudy (anonymous):

hello welcome to my question

OpenStudy (mathstudent55):

You are correct on the first question. Since the 750 in the equation is the initial amount she saved, when 750 becomes 700, it means the initial amount decreased from $750 to $700. For the second question: The point slope form of the equation of a line is \(y - y_1 = m(x - x_1) \) where m = slope, and the point is \((x_1, y_1) \) Now use the equation above, and just substitute the given values.

OpenStudy (anonymous):

so its b, correct?

OpenStudy (anonymous):

whoops

OpenStudy (mathstudent55):

Look carefully at the equation I wrote. Is x or y on the left side? Does the x term or the y term multiply m?

OpenStudy (anonymous):

so its D

OpenStudy (mathstudent55):

correct